# A Novel Spike Transformer Network for Depth Estimation from Event Cameras via Cross-modality Knowledge Distillation

### v1

@@ 2.1 Image-based and Event-based Monocular Depth Estimations
Depth estimation from images aims to measure the distance of each pixel relative to the camera. Monocular depth estimation, i.e. based on a single image, is a challenging but promising technology. It has the advantage of not requiring two images, which makes it more practical for applications where it is not possible to take a pair of images, such as mobile devices. Depending on the type of data used, we can divide the monocular depth estimation method into Image-based and Event-based methods [14]. The Image-based monocular depth estimation uses the information in an RGB image to estimate depth, while the event-based method uses the spike data generated by event camera. Image-based monocular depth estimation is more common than event-based monocular depth estimation because RGB images are easier to collect and process. Event cameras are a new type of sensors that output brightness changes in the form of an asynchronous "event" stream, rather than static images. This makes them well-suited for depth estimation in challenging conditions, such as low light and fast motion [33].
The latest developments in deep learning have made it possible to develop monocular depth estimation models that can achieved satisfactory accuracy and robustness[14, 34, 35].Similar to other deep learning models, these models typically consist of a generalized encoder that extracts abstract features from context information and a decoder that recovers depth information from the features. For RGB images, in [36], the author used ResNet-50 as an encoder and a novel up-sampling blocks as a decoder to estimate depth from a single RGB image. In [32], the authors utilised ViT instead of convolutional networks as the backbone for a depth estimation task. Experiments have found that transformer is able to provide finer and more globally consistent predictions than traditional convolutional networks. For event data, the authors [33] presented a new deep learning model called E2Depth that can estimate depth from event cameras with high accuracy. A fully convolutional neural network based on the U-Net architecture [37] were used in this work. In [38], a multiscale encoder was used to extract features from mixed-density event stacking and an upscaling decoder was used to predict the depth. The transformer structure was also used in event-based monocular depth estimation. In [39], EReFormer was proposed to estimate depth from event cameras with superior accuracy based on transformer.
@@ 4.1 Datasets
The first dataset is a synthetic dataset from [26], which is generated from DENSE dataset [33], including clear depth maps and intensity frames in 30 FPS under a variety of weather and illumination conditions. To obtain spike streams with high temporal resolution, the video is interpolated to generate intermediate RGB frames between adjacent 30-FPS frames. With absolute intensity information among RGB frames, each sensor pixel can continuously accumulate the light intensity with the spike generation mechanism, producing spike streams with a high temporal resolution (128$\times$30 FPS) that is 128 times of the video frame rate. The ‘spike’ version of DENSE dataset (namely DENSE spike) contains eight sequences, five for training, and three for evaluation. Each sequence consists of 999 samples, and each sample is a tuple of one RGB image, one depth map, and one spike stream. Each spike stream is simulated between two consecutive images, generating a binary sequence of 128 spike frames (with a size of 346 $\times$ 260 each) to depict the continuous process of dynamic scenes.
@@ 4.2 Experiment design
In this section, we evaluate the depth estimation performance and energy consumption of our SNN on the synthetic (DENSE) and real datasets (DSEC) and compare it with three competing dense prediction networks, namely U-Net [37], E2Depth [33] and Spike-T [26]. U-Net employs 2D convolutional layers as its encoder and focuses on spatial feature extraction, while E2Depth applies ConvLSTM layers that combine CNN and LSTM to capture the spatial and temporal features. The Spike-T employs transformer-based blocks to learn the spatio-temporal features simultaneously. These models therefore constitute our immediate and direct competitors.
@@ 5 Conclusion
Xin Zhang and Liangxiu Han conceived the experiment(s), Xin Zhang and Tam Sobeih conducted the experiment(s), Xin Zhang and Tam Sobeih analysed the results. All authors reviewed the manuscript.

### v2

@@ 2.1 Image-based and Event-based Monocular Depth Estimations
Depth estimation from images aims to measure the distance of each pixel relative to the camera. Monocular depth estimation, i.e. based on a single image, is a challenging but promising technology. It has the advantage of not requiring two images, which makes it more practical for applications where it is not possible to take a pair of images, such as mobile devices. Depending on the type of data used, we can divide the monocular depth estimation method into Image-based and Event-based methods [14]. The Image-based monocular depth estimation uses the information in an RGB image to estimate depth, while the event-based method uses the spike data generated by event camera. Image-based monocular depth estimation is more common than event-based monocular depth estimation because RGB images are easier to collect and process. Event cameras are a new type of sensors that output brightness changes in the form of an asynchronous "event" stream, rather than static images. This makes them well-suited for depth estimation in challenging conditions, such as low light and fast motion [Hidalgo-Carrió2020Learning].
The latest developments in deep learning have made it possible to develop monocular depth estimation models that can achieved satisfactory accuracy and robustness[14, 33, 34].Similar to other deep learning models, these models typically consist of a generalized encoder that extracts abstract features from context information and a decoder that recovers depth information from the features. For RGB images, in [35], the author used ResNet-50 as an encoder and a novel up-sampling blocks as a decoder to estimate depth from a single RGB image. In [32], the authors utilised ViT instead of convolutional networks as the backbone for a depth estimation task. Experiments have found that transformer is able to provide finer and more globally consistent predictions than traditional convolutional networks. For event data, the authors [Hidalgo-Carrió2020Learning] presented a new deep learning model called E2Depth that can estimate depth from event cameras with high accuracy. A fully convolutional neural network based on the U-Net architecture [36] were used in this work. In [37], a multiscale encoder was used to extract features from mixed-density event stacking and an upscaling decoder was used to predict the depth. The transformer structure was also used in event-based monocular depth estimation. In [38], EReFormer was proposed to estimate depth from event cameras with superior accuracy based on transformer.
@@ 4.1 Datasets
The first dataset is a synthetic dataset from [26], which is generated from DENSE dataset [Hidalgo-Carrió2020Learning], including clear depth maps and intensity frames in 30 FPS under a variety of weather and illumination conditions. To obtain spike streams with high temporal resolution, the video is interpolated to generate intermediate RGB frames between adjacent 30-FPS frames. With absolute intensity information among RGB frames, each sensor pixel can continuously accumulate the light intensity with the spike generation mechanism, producing spike streams with a high temporal resolution (128$\times$30 FPS) that is 128 times of the video frame rate. The ‘spike’ version of DENSE dataset (namely DENSE spike) contains eight sequences, five for training, and three for evaluation. Each sequence consists of 999 samples, and each sample is a tuple of one RGB image, one depth map, and one spike stream. Each spike stream is simulated between two consecutive images, generating a binary sequence of 128 spike frames (with a size of 346 $\times$ 260 each) to depict the continuous process of dynamic scenes.
@@ 4.2 Experiment design
In this section, we evaluate the depth estimation performance and energy consumption of our SNN on the synthetic (DENSE) and real datasets (DSEC) and compare it with three competing dense prediction networks, namely U-Net [36], E2Depth [Hidalgo-Carrió2020Learning] and Spike-T [26]. U-Net employs 2D convolutional layers as its encoder and focuses on spatial feature extraction, while E2Depth applies ConvLSTM layers that combine CNN and LSTM to capture the spatial and temporal features. The Spike-T employs transformer-based blocks to learn the spatio-temporal features simultaneously. These models therefore constitute our immediate and direct competitors.
@@ 4.4.2 Ablation study
5 Discussion
Transformer-based SNNs represent a novel form of SNNs that combines the transformer architecture with SNNs, providing great potential to break through the performance bottleneck of SNNs. Yao et al. [43], and Zhou et al[27], proposed two different Spike-Driven Self-Attention models. They utilised only mask and addition operations to avoid multiplication, which are efficient and have low computational energy consumption. Zhou et al [28], proposed Spikingformer , which modified the residual connection to purely event-driven, which is energy efficient while improving performance. However, the transformer based SNNs is still in its infancy in depth estimation[26]. There are two major issues: 1) the difficulty in training pure transformer based SNN models; 2) and the limited availability of paired depth data in event data to support transformer model training. Knowledge distillation provides a manner to train new or small models using well pretrained models. In this work, we propose a knowledge distillation method to bring ANN model knowledge into SNN.
This work introduces a novel spike transformer network for the monocular depth estimation task from event camera data. Our approach leverages the powerful spatial-temporal modeling capabilities of transformer architectures while adhering to the principles of spiking neural networks (SNNs) through the use of spike-driven residual learning and spike self-attention mechanisms.
Firstly, the proposed SNN backbone of spike transformers, driven purely by spikes, has demonstrated excellent capability and efficiency in extracting features from spike event data, achieving significant improvements in several evaluation metrics, resulting in satisfactory deep prediction results. Meanwhile, by eliminating the need for floating-point and integer-float multiplications, our approach significant energy savings while maintaining strong performance. This energy efficiency is a critical advantage for practical applications of SNNs[53, 54].
Moreover, the ablation study underscores the significance of the proposed fusion depth estimation head and knowledge distillation components. The fusion head, designed to fuse features from multiple transformer stages, proved helpful in preserving spatial and temporal information and recovering fine details in the depth predictions. Meanwhile, our proposed single-stage cross-modality knowledge distillation framework represents a novel and effective strategy for enhancing the performance of SNNs, even when faced with limited training data. By leveraging the knowledge encoded within DINOv2, a large-scale vision foundation model pre-trained on vast amounts of data, our SNN model can effectively capture relevant spatio-temporal features and mitigate the effects of noise and uncertainty inherent in spiking data, thereby improving its generalization capabilities. This approach effectively solves the overfitting problem faced by SNNs trained on small datasets, which seriously hinders the wide application of SNNs in various fields[5, 14].
The experimental results presented in this study provide compelling evidence of the efficacy of our proposed method. On both synthetic (DENSE) and real-world (DESC) event camera datasets, our model consistently outperformed existing state-of-the-art approaches, demonstrating substantial improvements in critical metrics such as absolute relative error and squared relative error. The visualization of depth predictions further highlights the model’s ability to capture intricate details, sharp edges, and contours, even in challenging low-light conditions.
While the results presented in this work are promising, further experiments and improvements are needed. For instance, deploying and evaluating the proposed SNN on neuromorphic computing hardware platform such as SpiNNaker[55], BrainScales[56] and TrueNorth[57]. As hardware acceleration for SNNs continues to advance, it will be crucial to assess the real-world performance and energy efficiency of our model in practical scenarios, such as autonomous vehicle navigation or robotic perception tasks.
It is also worth noting that while our work focused specifically on the task of depth estimation, the proposed spike transformer architecture and knowledge distillation framework could potentially be extended to other computer vision tasks. The inherent property of event camera data to capture high-rate temporal features provides an opportunity to utilise our proposed SNN in other dynamic scene tasks such as object tracking, motion estimation or event prediction. By leveraging the cross-modality knowledge distillation from the large foundation models and the powerful feature capture capabilities of transformer architecture, the proposed spike transformer has the ability to extract valuable features for the temporal evolution of complex environments, enabling more intelligent and reactive systems. This remains an attractive direction for future research.
@@ 6 Conclusion
X.Z. and LX.H. contributed to the conception and design of the study, while X.Z. and T.S. conducted the experiment(s). X.Z., LX.H., T.S., LH.H., and D.D. contributed to the analysis. LX.H. managed the project, acquired funding, and provided supervision. All authors contributed to the drafting and revision of the manuscript.